There is #30,000 in two accounts. One account has $4,000 more than the other. How much is in both accounts?
I know the anser is $17,000 and $13,000, nut I can't write it in an equation.

Answer by checkley77(12844)

X+(X+4,000)=30,000
X+X+4,000=30,000
2X=30,000-4,000
2X=26,000
X=26,000/2
X=$13,000 ANS
13,000+4,000=$17,000 ANS.